# Whether to trust: the ML leap of faith

### v1

@@ Abstract
Human trust is critical for trustworthy AI adoption. Trust is commonly understood as an attitude, but we cannot accurately measure this, nor manage it. We conflate trust in the overall system, ML, and ML’s component parts; so most users do not understand the leap of faith they take when they trust ML. Current efforts to build trust explain ML’s process, which can be hard for non-ML experts to comprehend because it is complex, and explanations are unrelated to their own (unarticulated) mental models. We propose an innovative way of directly building intrinsic trust in ML, by discerning and measuring the Leap of Faith (LoF)taken when a user trusts ML. Our LoF matrix identifies where an ML model aligns to a user’s own mental model. This match is rigorously yet practically identified by feeding the user’s data and objective function both into an ML model and an expert-validated rules-based AI model, a verified point of reference that can be tested a priori against a user’s own mental model. The LoF matrix visually contrasts the models’ outputs, so the remaining ML-reasoning leap of faith can be discerned. Our proposed trust metrics measure for the first time whether users demonstrate trust through their actions, and we link deserved trust to outcomes. Our contribution is significant because it enables empirical assessment and management of ML trust drivers, to support trustworthy ML adoption. Our approach is illustrated with a long-term high-stakes field study: a 3-month pilot of a sleep-improvement system with embedded AI .
@@ 1 Introduction
Lack of trust is often cited as a barrier to artificial intelligence (AI )adoption, especially in machine learning (ML). Trust has always been a critical enabler of new technology adoption: people tend to rely on automation they trust, and shun automation they distrust, in the real world [1] and the laboratory [2]. However, ML faces some unique challenges, which are discussed in the context of the psychology, management, and computer-science literatures.
@@
Trust is commonly understood as a unitary concept but, in practice, it is complex. Lee and Sees’ human-centred definition is most commonly accepted: the attitude that the technology will help us achieve our goals [3] in risky circumstances [4], i.e. in an uncertain and vulnerable situation [5]. Initial trust will vary for a technology because we each have a different general tendency to trust due to factors like our culture or personality (‘dispositional’ trust); and a different ability to deal with the situation and the technology (‘situational’ trust) [6]. Once we experience the specific technology, we can develop ‘learned’ trust. Most of the discussion of trust in the literature relates to cognitive or rational drivers of which the user is aware; but trust can also be emotional [7] and/ or have automatic or unconscious drivers, like impulses from learned associations and innate biases [8]. Lee and Sees’ ‘performance’/ ‘process’/ ‘purpose’framework [3] encapsulates most conscious drivers, but misses automatic ones and does not illuminate emotional ones. They argue potential for trust in technology is highest when: the trustee perceives ‘performance’ as strong; the trustee understands how the ‘process’ operates [9]; and the trustee believes its ‘purpose’ is aligned to their own goals. This framework is now used to reflect on the trust challenges faced by ML and tools to manage trust.
@@
ML should, in theory, garner higher trust when performance is better than alternatives. For example, a deep-learning model outperformed a team of 6 radiologists by a statistically-significant 11.5% in a 27,367-woman UK and US breast-cancer study, measured by relative area under the ROC curve [10]. However, if these technologists tested model output against what they expected given the inputs, they would disagree on outputs where the model outperformed them: false positives and negatives were 3.9-15.1% (UK-US) higher than the model. Despite this ML model’s performance, 4 years later, the NHS in the UK does not yet routinely use ML for breast-cancer screening, with trials still ongoing.
@@
The Explainable AI (XAI)movement has been motivated by a desire to improve our understanding of the ‘process’ of how ML models’ algorithms work [11]. However, because ML works by finding the best fit for mathematical models with large data sets [12], it can be hard to follow its reasoning process. Tools like feature importance or Shapley causal quasi values [13] reflect the factors that contribute to an ML model’s conclusion a posteriori [14], not its reasoning. If we understand a reasoning process and it matches our prior knowledge of sensible reasoning, we can directly build intrinsic trust; if not, only extrinsic trust is possible: we need external reassurance from an expert or an evaluation [4]. XAI’s efforts will only increase intrinsic trust if experts can explicitly compare the ML model’s reasoning to their own mental model [4]. This is problematic because current explanations do not always fit recipient mental models (their internal representation of how the world works [15]). For example, clinicians can think in terms of evidence-based mechanism of action [16]. Moreover, each human-in-the-loop may have a unique a priori mental model, which we only start to discover when ML outputs do not align with their implicit expectations. When time to resolve discrepancies is limited, intrinsic trust will not be established. Human systems of trust in high-stakes environments can rely on a trusted senior colleague pressure-testing a recommendation. Similarly, an expert human-in-the-loop can test whether a rules-based AI model reflects their mental model, and build intrinsic trust: they can interrogate the logic, testing whether it has been correctly implemented. When ML outperforms humans, especially on large data sets that are cognitively challenging to process, there should always be a mismatch between the model’s reasoning and expert priors, because new relationships are identified. It is thus not possible to directly build intrinsic trust in ML with experts through a collaborative process [17]. Trust can, at best, be extrinsic, relying on external assurances, so even experts require a leap of faith to act: from what they expect to what the model outputs. This leap can be challenging for experts with high self-confidence in their own abilities: they can ignore ML advice even though they know doing so lowers their performance [18].
However, understanding an ML model’s process requires comprehension of more than the algorithm: data and objective function are also critical. These could be understandable, which may require visualisation and education. An interpretable text feature space is: phrased in everyday language avoiding codes (‘readable’); worded so they are quick and simple to absorb (‘human-worded’); reliant on real-world concepts (‘understandable’) [19]. However, this approach does not include target variables, objective function, nor data visualisation, which can aid user when data sets are large.
@@
If a technology’s purpose is clearly communicated, it is generally trusted to do what it was built to do [3], although there could be concerns about how well it might do so. Because humans do not directly control how an ML algorithm reaches a conclusion, ML represents a fundamental shift in agency between humans and technology [20]; especially if outcomes are automated. If intent is not communicated, lack of agency could exacerbate concerns about purpose alignment, as could more general and/ or multi-agent ML: it will be harder to disentangle how its purpose relates to an individual’s own goals, or to trust it will make appropriate trade-offs, given other priorities.
Most of the literature on AI trust-management tools relates to rational drivers, namely performance, process and purpose. Glikson et al. explore tools with more emotional drivers: embodiment and immediacy behaviours [21]. Jacovi et al. suggest if we base our trust on such factors that are unrelated to model trustworthiness, our trust is undeserved [4]. However, a trustworthy system could ethically employ such techniques, e.g. if unconscious drivers preclude acceptance. It is harder to tell whether a system contains AI if the AI is embedded, or part of a larger system: the role of AI is less tangible and therefore harder to perceive, even if its presence is disclosed. Humans find embodied AI  – e.g. physical robots [22] – more tangible and thus easier to trust [21]. Trust in robots starts higher for the same level of model intelligence and deteriorates over time as performance is evaluated; whereas for embedded AI trust starts lower and builds over time [21]. However, underlying this insight is an issue with most trust measurement, especially embedded ML: it is trust in the system that is usually measured, not trust in the ML algorithm as distinct from data or objective function. Immediacy behaviours, like personalisation, which increase perceptions of interpersonal closeness [23], can improve perceived performance and purpose alignment independent of model trustworthiness. For example, psychological tailoring of advertising based on openness-to-experience or extroversion – assessed through digital footprint (e.g. Facebook likes) – increased 3.5 million users’ clicks by 40% and purchases by 50% [24]. Because AI is digital, it is easy to deploy users’ data for this purpose, and generative AI enables tailoring of tone. However, immediacy behaviours can backfire, especially when AI is embedded, reducing perceived decisional freedom and purpose alignment. Uber driver ‘algorithmic management’, with constant individual performance evaluation and feedback, increased driver awareness of the extent of ongoing monitoring, making them resentful of being micro-managed by an algorithm whose purpose was perceived not as not aligned to their goals [25].
We lack a hierarchy of trust drivers and tools to support trustworthy ML adoption. A critical issue is our inability to directly measure attitudinal trust, so we cannot objectively assess what supports it. We can only measure declared, demonstrated and, with our proposed metric, deserved trust, which we now discuss; see Figure 1 for a diagram of how these fit together.
Figure 1: Declared, demonstrated and deserved trust, and how they differ from attitudinal trust. All human-centred concepts are shown in yellow; technology ones in blue; combined in green. Icons by Angriawan Ditya Zulkamain, Arif Hariyanto, Karen Tyler and PEBIAN from thenounproject.com.
Declared trust, what people say they trust, has been the primary focus for measurement [26]. This is derived from attitudinal trust because it only reveals what the respondent is conscious of and is prepared to disclose, given social norms. It will therefore miss automatic drivers of which the respondent is unaware, like biases. Declared trust is analysed using subjective data from interviews and surveys, using different scales so it is hard to learn across studies [27]. Crucial trust drivers may be missing because scales are researcher defined.
Demonstrated trust is a behaviour: we take a risk and act, displaying ‘reliance’ [3]. Demonstrated trust reflects both conscious and automatic drivers but it is not the same as attitudinal trust, e.g. a user may trust but not act because they lack behavioural control, or fear being judged by society; or they might set an intention to act, but not follow-through because of constraints encountered [28]. Demonstrated trust can be objectively measured, e.g. Weight on Advice [18] or percentage of recommendations relied upon [29]. However, these metrics fail to reflect action follow-through, which is necessary when there are constraints to be overcome. Furthermore, there is little exploration of trust-management-tool implications, e.g. there could be a difference between what boosts declared trust and follow-through to action. The relationship between declared trust and action appears strongest when there is high cognitive complexity, e.g. complex automation and novel situations [3]. Reliance can also be stronger when the individual has high decisional freedom as to subsequent action and is able to compare the technology’s recommendation to the human alternative [6].
Warranted or deserved trust is when the risk was worth it because the technology itself was trustworthy [3]. Trustworthiness is a property of technology, not an individual, so it is not necessary for attitudinal or demonstrated trust [3]. Trustworthiness is ideally calibrated to trust so technology is not misused, disused or abused, reducing safety risks [30]. There is substantial discussion on trustworthiness measurement, e.g. the EU criteria include human agency and oversight; transparency; diversity, non-discrimination, fairness; as well as prevention of harm: technical robustness and safety; privacy and data governance; accountability; and societal and environmental well-being [31]. However, the literature on deserved-trust measurement is limited. Jacovi et al. suggest model trustworthiness should be manipulated so it is no longer calibrated with attitudinal trust, to see how much attitudinal trust changes [4]. This is a conceptual approach that is challenging to put into practice in the real world in a verifiable, controlled, reproducible, comparable manner. This approach fails to assess what we really want to know: did trusting the recommendation deliver better outcomes? While a model might be technically robust, it might not take into account factors that would influence an individual’s perception of behavioural control or situation-specific constraints. In anything involving behaviour change, deserved trust is derived from demonstrated trust, not attitudinal, because a recommendation needs to be acted on for impact to be assessed. The individual’s experience of the specific situation and technology, learning whether trust was deserved or not, then translates back to attitudinal trust.
@@
Most users do not therefore understand the leap of faith they take when they trust ML. Current efforts to build trust fail because explanations are unrelated to domain experts’ (unarticulated) mental models. We propose a more practical way of building trust, that shortens the leap. We now illustrate this using data from a long-term high-stakes field study: a 3-month pilot for a sleep-improvement system with embedded AI . We briefly describe our study methods in Section 2, before we propose our novel architecture in Section 3. Section 4 evaluates the contribution of this architecture and its limitations.
@@ 2 Methods
The study was high stakes because it impacted health [32], and the scale of changes were substantial. Most participants were short sleepers – i.e. they slept less than 7 hours a night – which is associated with poor health outcomes, e.g. 1.2-1.4x higher Alzheimer’s disease risk [33]. Immune response suffers, so short sleepers are 5x more susceptible to infections like the common cold [34]. Reaction times [35] decline, similar to alcohol intoxication [36]. The main treatment for those who are not chronic insomniacs is ‘sleep hygiene’: up to 16 behavioural and environmental suggestions (e.g. avoid bright light in the evening). While these interventions have a medium effect in healthy-adult sleep trials [37], research has been subjective and failed to take into account individual sensitivity, even though it varies greatly: e.g. one person’s sensitivity to evening light can be 40x that of another person [38]. Slow-Wave Sleep (SWS)declines as we age [39] and clears beta amyloid plaques, linked to Alzheimer’s disease, from the brain [40]. There has been limited research into the impact of these practices on healthy-adult SWS. The situation was thus uncertain and participants vulnerable, so they took a risk when they followed a recommendation.
@@
Figure 2: Socio-technical sleep-improvement system with embedded parallel neuro-symbolic AI.
Our models were designed to be trustworthy using the EU criteria [31]: human agency and oversight, and transparency are central our socio-technical-system design. We implemented the other principles as follows: a) Models were run on each individual’s own data. b) Participants were informed of model technical robustness and safety: we informed them that rules-based model recommendations were sleep-expert approved and of their own ML model accuracy: 90% (87-94% range, 2.4% SD); and 80% for SWS 80% (75-84% range, 3.3% SD). c) Privacy and data standards were specified by the University Ethics Committee (PREC reference number 22 003, 5th September 2022). These standards were a key selection criteria for smart devices, requiring the team to create their own environmental monitor because no third-party solution was identified that delivered sufficient granularity and met these standards. d) The University Ethics Committee holds the researcher accountable for model trustworthiness and upholding standards, and each participant was accountable for their own data, objective function, and choices. e) The key societal challenge was ensuring participant sleep anxiety did not increase as a consequence of the study. Sleep anxiety was measured before and after the study, and it (statistically insignificantly) declined. Severe anxiety or depression were screened out, through GAD7 and PHQ9 questionnaires. Sleep end points were continually monitored, with an 85% floor on Sleep Efficiency (SE), the percentage of time participants were asleep while in bed.

### v2

@@ Abstract
Human trust is a prerequisite to trustworthy AI adoption, yet trust remains poorly understood. Trust is often described as an attitude, but attitudes cannot be reliably measured or managed. Additionally, humans frequently conflate trust in an AI system, its machine learning (ML) technology, and its other component parts. Without fully understanding the ‘leap of faith’ involved in trusting ML, users cannot develop intrinsic trust in these systems. A common approach to building trust is to explain a ML model’s reasoning process. However, such explanations often fail to resonate with non-experts due to the inherent complexity of ML systems and explanations are disconnected from users’ own (unarticulated) mental models. This work puts forward an innovative way of directly building intrinsic trust in ML, by discerning and measuring the Leap of Faith (LoF)taken when a user decides to rely on ML. The LoF matrix captures the alignment between an ML model and a human expert’s mental model. This match is rigorously and practically identified by feeding the user’s data and objective function into both an ML agent and an expert-validated rules-based agent—a verified point of reference that can be tested a priori against a user’s own mental model. This represents a new class of neuro-symbolic architecture. The LoF matrix visually contrasts the different agent outputs, revealing to the user the distance that constitutes the leap of faith between the rules-based and ML agents. For the first time, we propose trust metrics that evaluate whether users demonstrate trust through their actions rather than self-reported intent and whether such trust is deserved based on outcomes. The significance of the contribution is that it enables empirical assessment and management of ML trust drivers, to support trustworthy ML adoption. The approach is illustrated through a long-term high-stakes field study: a 3-month pilot of a multi-agent sleep-improvement system.
@@ 1 Introduction
Lack of trust is often cited as a barrier to artificial intelligence (AI )adoption, especially for machine learning (ML). Trust has always been a critical enabler of new technology adoption: people tend to rely on automation they trust, and shun automation they distrust, in the real world [1] and the laboratory [2]. However, ML faces some unique challenges, which are discussed in the context of the psychology, management, and computer-science literatures.
@@
Trust is commonly understood as a unitary concept but, in practice, it is complex. Lee and Sees’ human-centred definition is widely accepted: the attitude that the technology will help us achieve our goals [3] in risky circumstances [4], i.e. in an uncertain and vulnerable situation [5]. Initial trust will vary for a technology because we each have a different general tendency to trust due to factors like our culture or personality (‘dispositional’ trust); and a different ability to deal with the situation and the technology (‘situational’ trust) [6]. Once we experience a specific technology, we can develop ‘learned’ trust. Most of the discussion of trust in the literature relates to cognitive or rational drivers of which the user is aware. Lee and Sees’ ‘performance’/ ‘process’/ ‘purpose’ framework [3] encapsulates most of these conscious drivers. They argue potential for trust in technology is highest when: the trustee perceives ‘performance’ as strong; the trustee understands how the ‘process’ operates [7]; and the trustee believes its ‘purpose’ is aligned to their own goals. This framework is now used to reflect on the trust challenges faced by ML.
@@
ML should, in theory, garner higher trust when performance is better than alternatives. For example, a deep-learning model outperformed a team of 6 radiologists by a statistically-significant 11.5% in a 27,367-woman UK and US breast-cancer study, measured by relative area under the ROC curve [8]. However, if these technologists tested model output against what they expected given the inputs, they would disagree on outputs where the model outperformed them: false positives and negatives were 3.9-15.1% (UK-US) higher than the model.
@@
Explainable AI (XAI)research motivated by a desire to improve our understanding of the ‘process’ of how ML models’ algorithms work [9]. If we understand a reasoning process and it matches our prior knowledge of sensible reasoning, we can directly build intrinsic trust; if not, only extrinsic trust is possible: we need external reassurance from an expert or an evaluation [4]. XAI’s efforts will only increase intrinsic trust if humans can explicitly compare the ML model’s reasoning to their own mental model [4] and their internal representation of how the world works [10]. This is problematic because a given explanation does not always fit recipient mental models. Because ML works by finding the best fit for mathematical models with large data sets [11], it can be challenging for people to follow its reasoning process because it is so different from their own. Tools like feature importance or Shapley causal quasi-values [12] reflect the factors that contributed to an ML model’s conclusion a posteriori [13], but not its reasoning. For example, clinicians can think in terms of an evidence-based mechanism of action [14]. Each human-in-the-loop may have a unique a priori mental model, which we only start to discover when ML outputs do not align with their implicit expectations. When ML outperforms humans, there should always be a mismatch between the model’s reasoning and expert priors, because new relationships are identified. Trust can thus, at best, be extrinsic, relying on external assurances, so even domain experts require a leap of faith to act: from what they expect to what the model outputs. However, understanding a ML model’s ‘process’ requires comprehension of more than algorithmic reasoning: data is also critical and could be understandable, even if the data set is large. Zytek et al. propose that an interpretable text feature space is: phrased in everyday language avoiding codes (‘readable’); worded so concepts are quick and simple to absorb (‘human-worded’); and reliant on real-world concepts (‘understandable’) [15].
@@
If a technology’s purpose is clearly communicated, it is generally trusted to do what it was built to do [3], although there could be concerns about its performance. Because humans do not directly control how an ML algorithm reaches a conclusion, ML represents a fundamental shift in agency between humans and technology [16]; especially if workflows are automated, and made up of agents that perceive and act [17]. If intent is not communicated, lack of agency could exacerbate concerns about purpose alignment, especially if a system has multiple agents: it will be harder to disentangle how its purpose relates to an individual’s own goals, or to trust it will make appropriate trade-offs, given other priorities.
Trust can also be emotional [18] and/ or have automatic or unconscious drivers, like impulses from learned associations and innate biases [19]. Jacovi et al. suggest if we base our trust on such factors that are unrelated to model trustworthiness, our trust is undeserved [4]. However, a trustworthy system could ethically employ such techniques e.g., if unconscious drivers preclude acceptance. Glikson et al. explore some relevant tools: embodiment and immediacy behaviours [20]. Humans find embodied AI  – e.g. physical robots [21] – more tangible and thus easier to trust [20]. It is harder to tell whether a system contains AI if the AI is embedded, or is part of a larger system: the role of AI is less tangible and therefore harder to perceive, even if its presence is disclosed. Agents are digitally, not physically, embodied so sit between the two. Immediacy behaviours, such as personalisation, increase perceptions of interpersonal closeness [22] so can improve perceived performance and purpose alignment.
@@
Following the above brief presentation of aspects of trust, we propose a practical way of measuring and managing trust in ML, that discerns the leap of faith, and measures both demonstrated and deserved trust, enabling trust to be managed over time. We illustrate this using data from a long-term high-stakes field study: a 3-month pilot for a sleep-improvement system with AI agents. We first review related work in Section 2, then briefly describe our study methods in Section 3 to provide the context for the novel architecture we propose in Section 4. Section 5 evaluates the contribution of the architecture and its limitations.
2 Related work
The literature lacks a hierarchy of tools to support trustworthy ML adoption. Our inability to directly measure attitudinal trust means that we cannot objectively assess what drives it. However, what matters most for AI adoption is whether people rely on AI recommendations, and whether this trust was deserved. These concepts are depicted in Figure 1 and will now be explored. We conclude that new metrics are required to measure demonstrated and deserved trust.
Figure 1: Trust and trustworthiness definitions. All human-centred concepts are shown in yellow; technology ones in blue; combined in green. Icons by Angriawan Ditya Zulkamain, Arif Hariyanto, Karen Tyler and PEBIAN from thenounproject.com.
Declared trust
What is routinely measured, and labeled as attitudinal trust, is declared trust or what people say they trust [23]. It is wrong to label this attitudinal trust, because it only reveals what the respondent is conscious of and is prepared to disclose, given social norms, so it will miss automatic drivers of which the user is unaware, like biases. Declared trust is measured subjectively in interviews and surveys using scales that are different and researcher-defined, so it is hard to learn across studies [24] and key trust drivers may be missed. Declared trust in robots starts higher for the same level of model intelligence and deteriorates over time as performance is evaluated; whereas for embedded AI trust starts lower and builds over time [20]. Underlying this insight is an issue with most trust measurement, especially for embedded ML: it is trust in the system that is measured, not trust in the ML algorithm as distinct from data or objective function.
Demonstrated trust
Demonstrated trust is an action or a behaviour: we take a risk and act, displaying ‘reliance’ [3]. Demonstrated trust reflects both conscious and automatic drivers but it is not the same as attitudinal trust  – e.g. a user may trust but not act because they lack behavioural control, or fear being judged by society; or they might set an intention to act, but not fulfill this intention (or ‘follow through’) because of constraints encountered [25]. Demonstrated trust can be objectively measured. For example, percentage of recommendations relied upon [26] or Weight on Advice, which measures the extent to which participants adjust their decisions based on the advice provided by an algorithm or a human [27]. However, these metrics only measure intention and fail to reflect action follow-through, which is critical for adoption, especially if there are multiple constraints to overcome. The relationship between declared trust and follow-through to action may be strongest when there is high cognitive complexity  – e.g. complex automation and novel situations [3]. Reliance can also be stronger when an individual has high decisional freedom as to subsequent action and is able to compare the technology’s recommendation to the human alternative [6].
Deserved trust
Warranted, or deserved trust, is when the risk was worth it because the technology itself was trustworthy for the use-case [3]. Trustworthiness is a property of technology, not a human, so is independent of trust [3]: humans can distrust and therefore fail to adopt trustworthy AI ; humans can trust and adopt untrustworthy AI . Trustworthiness is ideally calibrated to trust so technology is not misused, disused or abused, reducing safety risks [28].
There is substantial discussion on measuring model trustworthiness  – e.g. the EU criteria include human agency and oversight; transparency; diversity, non-discrimination, fairness; as well as prevention of harm: technical robustness and safety; privacy and data governance; accountability; and societal and environmental well-being [29].
However, the literature on measuring deserved trust is limited – i.e. was the individual right to act upon their trust in the model? Jacovi et al. suggest one approach is to manipulate model trustworthiness so it is no longer calibrated with attitudinal trust, and see how much attitudinal trust changes [4]. This is a conceptual approach that is challenging to put into practice in the real world in a verifiable, controlled, reproducible, comparable manner. Furthermore, it suggests trust changes in a linear manner which may not be true (e.g. it could be step-wise). This approach also fails to assess what we really want to know: did trusting the recommendation deliver better outcomes? Demonstrated trust is key, not attitudinal, because a recommendation needs to be acted on for impact to be assessed. The individual’s experience in a specific use-case will provide learning as to whether trust was deserved or not, which should then translate back into attitudinal trust and, over time, declared and demonstrated trust. While a model might be technically robust, it might not have taken into account factors that would influence action – e.g. an individual’s perception of behavioural control or situation-specific constraints. There has been very limited exploration of outcomes in the literature. When mortality outcomes for clinician decisions in treating sepsis in intensive care were assessed , 90-day mortality was found to be lowest when the doctors’ dosage was in line with an RL model’s recommendations [30], indicating this model was deserving of trust. The issue with this experiment is that it is that the researchers did not have the option to follow a ML recommendation rather than their own, so the comparison was done ex-post. The value of following the ML recommendation is therefore not directly evaluated – it is modeled, using a high-confidence off-policy evaluation method. Other examples in this space simulate clinical environments [31], rather than providing decision makers with decision-making tools in the real world and assessing the value when decision makers follow an ML recommendation.
@@ 3 Methods
The sleep-improvement study was high stakes because it impacted health [32], and the scale of required changes were substantial. Most participants were short sleepers – i.e. their Total Sleep Time (TST)was less than 7 hours a night  – which is associated with poor health outcomes, e.g. 1.2-1.4x higher Alzheimer’s disease risk [33]; lower immune response, so short sleepers are 5x more susceptible to infections like the common cold [34]; slower reaction times [35], similar to alcohol intoxication [36] leading to accidents – e.g. 1.8x higher in the workplace [37]. The main treatment for those who are not chronic insomniacs is ‘sleep hygiene’ (SH): up to 16 behavioural and environmental suggestions (e.g. avoid light in the evening). While SH interventions have a medium effect in healthy-adult sleep trials [38], research has been subjective and failed to take into account individual sensitivity, even though it varies greatly: e.g. one person’s sensitivity to evening light can be 40x that of another person [39]. Slow-Wave Sleep (SWS)clears beta amyloid plaques, linked to Alzheimer’s disease, from the brain [40]. There has been limited research into the impact of SH practices on healthy-adult SWS. The situation was thus uncertain, so participants took a risk when they followed a recommendation.
@@
Figure 2: Socio-technical sleep-improvement system with parallel neuro-symbolic AI.
Our models were designed to be trustworthy using the EU criteria [29]: human agency and oversight, and transparency are central our socio-technical-system design. We implemented the other principles as follows: 1. Models were run on each individual’s own data. 2. Participants were informed of model technical robustness and safety: we informed them that rules-based model’s recommendations were sleep-expert approved and of their own ML model accuracy: 89% for TST (81-92% range, 3.5% SD); and 81% for SWS (75-86%, 4.5% SD). 3. Privacy and data standards were specified by the University Ethics Committee (PREC reference number 22 003, 5th September 2022). These standards were a key selection criteria for smart devices, requiring the team to create their own environmental monitor because no third-party solution was identified that delivered sufficient granularity and met these standards. 4. The University Ethics Committee holds the researcher accountable for model trustworthiness and upholding standards, and each participant was accountable for their own data, objective function, and choices. 5. The key societal challenge was ensuring participant sleep anxiety did not increase as a consequence of the study. Sleep anxiety was measured before and after the study, and it (statistically insignificantly) declined. Severe anxiety or depression were screened out, through GAD7 and PHQ9 questionnaires. Sleep end points were continually monitored, with an 85% floor on Sleep Efficiency, the percentage of time participants were asleep while in bed.